Write the Slope Intercept form of an equation for the line that passes through (4,0) and is parallel to the graph of 3y-6x=4
Answer by stanbon(75887)

Write the Slope Intercept form of an equation for the line that passes through (4,0) and is parallel to the graph of 3y-6x=4
--------------------------
Find the slope of the given equation:
y = 2x+(4/3)
slope = 2
-------
Find the equation of the line with slope = 2 and y = 0 when x = 4.
---------------------
Form y = mx + b
0 = 2*4 + b
b = -8
-------
Equation you wnat:
y = 2x-8
==============
Cheers,
Stan H.
=============